# Fast and low energy approximate full adder based on FELIX logicthanks: Preprint Submitted to Journal of Engineering

### v2

@@ Abstract
Nowadays, much data must be processed and moved between processing and memory units. New technologies and architectures have emerged to improve system performance and overcome the memory bottleneck. The memristor is a technology with both computing and memory capabilities. In-Memory Computing (IMC) can be performed by applying memristors to stateful design methods. The Fast and Energy-Efficient Logic in Memory (FELIX) logic is a stateful implementation logic. The way computations are performed can be changed to improve performance. Approximate design methods can be applied in error-resilient applications. In this paper, an approximate full adder circuit with exact $C_{out}$ and approximate $Sum$ outputs has been proposed using the FELIX design method for IMC in two different implementation approaches. The memristor count in the proposed FELIX-based Approximate Full Adder (FAFA) in the two proposed implementation approaches is improved by 14.28% and 28.57%, energy consumption is improved by 73.735% and 81.754%, and the number of computational steps in both approaches is improved by 66.66% compared to the exact FELIX-based full adder. Two different scenarios are considered for evaluating the FAFA. The results of error analysis and evaluations of the FAFA in three different image processing applications confirmed that FAFA has high accuracy and acceptable performance.
@@
Memristor, Approximate Computing, Approximate Full Adder, In-Memory Computing, FELIX, Image Processing.
@@ 1 Introduction
Recently, numerous efforts have been undertaken to address the Von-Neumann bottleneck or memory wall. This bottleneck arises from the significant energy consumption and delays in transferring data between memory and processing units [1]. Most of the computational energy comes from the energy required to move data between these units because the energy to access memory increases dramatically along the memory hierarchy (from cache to off-chip memory) [2, 3]. Pedram et al. [4] have stated that the energy consumption of a 16-bit adder in 45 nm Complementary Metal-Oxide Semiconductor (CMOS) technology is 3556 times less than the energy needed to access Dynamic Random Access Memory (DRAM). Also, the DRAM access time is 25 times higher than the 32-bit adder delay in CMOS technology [2]. This problem significantly reduces system performance in applications where the volume of data is large and calculations are complex (e.g., machine learning, neural networks, video, and image processing). The physical proximity of memory and processing units was proposed to reduce energy consumption and data movement delay [5]. However, this solution is only able to eliminate data movement partially. On the other hand, reducing the size of MOS Field Effect Transistors (MOSFETs) has faced limitations. These limitations include reduced reliability and gate control, increased subthreshold leakage currents, leakage power, and a cost wall [6, 7, 8]. Consequently, alternative technologies are required to overcome the above problems. Research has been carried out in the past decades, which resulted in a shift towards emerging technologies such as Carbon Nanotube FET (CNFET) [3], Quantum dot Cellular Automata (QCA), Ferroelectric FET (FEFET), nanomagnet logic, and memristors [3, 9].
@@
Adders and multipliers are applied in almost every arithmetic calculation in a computer system [13]. Therefore, because the performance of the systems is influenced by the wide applicability of these units in numerous applications and the high volume of data required for processing, an approximate full adder circuit for IMC has been proposed in this paper. The proposed FELIX-based Approximate Full Adder (FAFA) can be used well in error-resilient applications since the erroneous bit is not propagated from the Least Significant Bits (LSBs) structure to the Most Significant Bits (MSBs). The FAFA has been designed using the FELIX logic method, which is compatible with the structure of crossbar arrays. The key contribution of this paper is proposing a high-speed approximate full adder implementation algorithm (FAFA) with low energy consumption applicable in IMC alongside presenting an approximate truth table with exact $C_{out}$. Optimizing the execution algorithm of the $1^{st}$ implementation approach (FAFA1), following the basic principles of the FELIX method, and presenting a new optimized implementation algorithm (FAFA2) to improve the circuit-level evaluation criteria significantly was the main novelty of the authors in this paper. We calculate circuit evaluation metrics improvements of the FAFA1 and FAFA2 compared to the exact FELIX-based full adder and different exact and approximate IMPLY-based full adders. In order to ensure the output’s accuracy, error analysis metrics in different scenarios are assessed. Then, we apply the FAFA circuit in image processing applications and evaluate it according to valid measurements to analyze the quality of images.
@@ 2.1 Memristors
In 1971, Leon Chua proposed the memristor as the fourth fundamental component linking flux and charge [23]. It wasn’t until 2008 that Stanley Williams’ team at HP Labs successfully constructed the first memristor [24]. Following this, researchers have investigated and utilized memristors in various fields. A memristor is a switchable resistor with two electrodes and an insulating active layer between them. Figure 1 shows a memristor [25]. The resistance of a memristor is influenced by the previous voltage or current that has been applied to it. The value of this resistor can change between two values, $R_{on}$ (low resistance state, logic ‘1’) and $R_{off}$ (high resistance state, logic ‘0’), based on the voltage or current applied to it. Suppose the current is applied from the cathode direction (as shown in Figure 1) or the voltage applied to the memristor is greater than the threshold voltage of the memristor. In that case, the memristor’s resistance decreases and reaches $R_{on}$. On the other hand, when current is applied from the anode direction or a voltage below the negative threshold is applied, the resistance increases, causing the memristor to switch to its $R_{off}$ state.
@@ 2.2 FELIX Design Method
Among the memristor-based design methods, one of the inputs is rewritten in the IMPLY method. In the MAGIC method, only the NOR function is implemented in the crossbar array, and all other functions are obtained using NOR, which leads to unnecessary delay overhead. The MRL method is unsuitable for IMC; the FELIX design method is applied in this paper. The FELIX design method enables the activation of IMC and many Boolean functions such as OR, NOT, NOR, Minority (MIN), and NAND in only one cycle and the Majority (MAJ) and AND functions in two cycles without any additional memristor, and the XOR function in two cycles and an additional memristor to run in the crossbar array [16]. While implementing OR in MAGIC requires two NOR cycles, and implementing MAJ and AND requires four cycles in this design method [16]. Like the MAGIC design method, the FELIX design method preserves inputs, and thus, the input data set can be reused for multiple operations [30]. In the FELIX design method, to switch the memristor from $R_{off}$ to $R_{on}$, $V_{pn}>|v_{on}|$ must be established, and $V_{np}>|v_{off}|$ must be established for switching from $R_{on}$ to $R_{off}$. $|v_{on}|$ and $|v_{off}|$ are the memristor’s threshold voltages, and $V_{pn}$ ($V_{np}$) is the voltage difference between the $p$ ($n$) and $n$ ($p$) terminals [16]. In FELIX, the execution voltage, $V_{0}$, is a key factor. A specific $V_{0}$ voltage for each logic function should be applied to operate. In other words, the applied voltage determines the type of operation. Figure 2 shows the implementation of 3-input NOR, 3-input NAND, and MIN gates [16]. As shown in Figure 2, these three gates have the same implementation, and the applied voltage value, $V_{0}$, determines the type of operation. These operations require only one cycle. For example, implementing a 3-input NOR gate in a memristor-based crossbar array is described. In the first step, the output memristor (labeled out in Figure 2) is initialized to $R_{on}$. In the next step, $V_{0}$ is applied from the cathode direction (terminal $p$) of the inputs, and the cathode side of the output memristor is connected to the ground.
@@ 2.3 Memristor-Based Adders
Several methods have been proposed in the literature for designing and implementing full adder cells using memristors. These methods include IMPLY, MAGIC, FELIX, reconfigurable NAND/AND, and MRL. The authors in [21, 31] introduced IMPLY-based exact full adders using serial architecture. Considering the initialization of work memristors, the full adder in [21] needs five memristors. It computes the outputs in 22 computational steps, and the full adder proposed in [31] needs five memristors and 23 computational steps to compute $Sum$ and $C_{out}$. In [32], an exact Carry Select Adder (CSA) and an exact Conditional Carry Adder (CCA) are introduced using parallel architecture and with the IMPLY method. The proposed 4-bit CSA and CCA require 44 and 49 memristors, respectively, and take 39 and 41 computational steps, respectively [32]. Compared to serial architecture, parallel architecture has the advantage of being faster. Still, it has disadvantages, such as the need for switches, more memristors, high hardware complexity, and a lack of full compatibility with crossbar arrays [1].
In [16], in addition to introducing the FELIX method and implementing basic logic gates, an exact full adder cell for IMC using XOR and MAJ functions is proposed. In [16], two XOR functions are required to generate $Sum$ output, which takes four cycles and requires two additional memristors. The $C_{out}$ output is also generated after two cycles and requires two additional memristors. It should be mentioned that the initialization of memristors also requires a computational cycle, so the number of cycles increases to seven cycles.
Implementing a full adder using the MAGIC method requires twelve NOR cycles and twelve additional memristors compared to six cycles and four additional memristors in FELIX [33]. The operations related to the exact full adder proposed in [16] and its memristive circuit are illustrated in Table 1 and Figure [3], respectively. In Table 1 and Figure 3, $A_{in}$, $B_{in}$, and $C_{in}$ are the inputs of the full adder, and $W_{1}$ to $W_{4}$ are the four additional memristors needed to perform the addition operation.
@@
Figure 3: Circuit-level design of the FELIX-based full cell in [16].
@@
The authors in [1, 22] proposed approximate full adders in the serial architecture using the IMPLY method. The design proposed in [1] requires four memristors and seven computational steps to calculate the outputs, while the best approximate full adder proposed in [22] needs four memristors and eight computational steps.
@@
[16] 0 0 0 0 0 7 6 - ✓ (FELIX) Exact
[33] 0 0 0 0 0 15 12 - ✓ (MAGIC) Exact
@@ 3 Proposed Memristive Approximate Full Adder Based On FELIX
In the second approach, the proposed circuit involves five memristors (three input memristors, $A_{in}$, $B_{in}$, and $C_{in}$, and two additional memristors, $W_{1}$ and $W_{2}$). The whole operation is performed in two cycles: one to calculate the MIN function and one to perform the MAGIC NOT operation. In both implementation approaches, outputs are calculated in three cycles if the initialization of memristors is considered. The implementation algorithms of the FAFA1 and FAFA2 are written in Tables 4 and 5, respectively, and their circuit-level designs are shown in Figures 4(a) and 4(b).
@@ 4.1 Circuit-Level Simulation And Analysis
The LTSPICE simulator verified the correctness of the outputs of all eight possible input states of the FAFA1 and FAFA2. Circuit-level simulation results showed that the proposed algorithm produces outputs corresponding to the proposed truth table in all states. The outputs of the proposed circuit in two input states, $A_{in}B_{in}C_{in}$=“111” and $A_{in}B_{in}C_{in}$=“010”, in the first and second approaches are shown in Figures 5(a) and 5(b) and 6(a) and 6(b), respectively.
@@
Table 7 lists the circuit characteristics of the stateful memristive exact and approximate full adders proposed in [1, 21, 22], which are implemented using the IMPLY design method in serial architecture. This table compares only the number of required memristors and the computational steps of the proposed approximate full adders (FAFA1 and FAFA2) with the exact and approximate IMPLY-based full adders. Due to the difference in the memristor model used in the circuit level simulation of this paper with the proposed circuits in [1, 21, 22], it is not possible to directly compare the energy consumption of these circuits with each other. The number of memristors applied in the IMPLY-based exact and approximate full adders proposed in [1, 21, 22] is at most 33.33% less than the number of memristors applied in the proposed circuits here. However, it should be noted that the IMPLY method is destructive, and the input values stored in the input memristors are erased after the implementation algorithm is executed. Therefore, if input data needs to be stored, a copy operation (two NOT operations for each input bit) must be performed using work memristors, significantly increasing the number of computational steps and the number of memristors. The FELIX method is non-destructive, and there is no need to perform a copy operation to preserve the data after executing the proposed circuit implementation algorithm. Proposed approximate full adders have improved the number of computational steps by 66.66%-90% compared to the proposed circuits in [1, 21, 22]. The mentioned percentages are reported without considering the computational steps needed to store the input data in the IMPLY-based exact and approximate full adders [1, 21, 22].
Figure 5: The output waveform of the FAFA1 in the states: (a) $A_{in}B_{in}C_{in}=``010$", and (b) $A_{in}B_{in}C_{in}=``111$".
Figure 6: The output waveform of the FAFA2 in the states: (a) $A_{in}B_{in}C_{in}=``010$", and (b) $A_{in}B_{in}C_{in}=``111$".
Table 7: Circuit-level criteria comparison.
@@
Exact IMPLY [21] 5 28.57% 20 -233.33% NC [^1] Not Calculated NC
SIAFA1 [22] 4 42.85% 7 -16.66% NC NC
SIAFA2 [22] 5 28.57% 8 -33.33% NC NC
SIAFA3 [22] 4 42.85% 7 -16.66% NC NC
SIAFA4 [22] 4 42.85% 7 -16.66% NC NC
SAFAN [1] 4 42.85% 6 0% NC NC
@@
For a comprehensive analysis of the FAFA, the 8-bit approximate Ripple Carry Adder (RCA) has been evaluated in two different scenarios. The proposed 8-bit approximate adder was divided into two exact and approximate parts in both scenarios. In the exact part of the $1^{st}$ scenario’s structure, the exact full adder was applied to the four MSBs. The FAFA was applied in the LSBs in this scenario. In the $2^{nd}$ scenario, the exact section contains three exact full adders in the MSBs, and the other five are approximate ones (LSBs). In both scenarios, the exact full adder proposed in [16] is used for the MSBs. The $1^{st}$ scenario’s structure is drawn in Figure 7. The memristors’ count, cycles, and the energy consumption of the proposed approximate 8-bit RCA in the FAFA1, FAFA2, and in both scenarios are compared with the exact 8-bit RCA built using the exact full adder proposed in [16] in Table 8.
@@
Table 8: Comparison of the number of memristors, the number of cycles, and the energy consumption of the proposed 8-bit approximate RCA with the 8-bit exact RCA proposed in [16].
@@
SIAFA1 [22] 19 45.71% 120 -114.28% NC NC
SIAFA2 [22] 19 45.71% 128 -128.57% NC NC
SIAFA3 [22] 19 45.71% 120 -114.28% NC NC
SIAFA4 [22] 19 45.71% 120 -114.28% NC NC
SAFAN [1] 19 45.71% 116 -107.14% NC NC
@@
SIAFA1 [22] 19 45.71% 106 -89.28% NC NC
SIAFA2 [22] 19 45.71% 116 -107.14% NC NC
SIAFA3 [22] 19 45.71% 106 -89.28% NC NC
SIAFA4 [22] 19 45.71% 106 -89.28% NC NC
SAFAN [1] 19 45.71% 101 -80.35% NC NC

### v3

# Fast and low-energy approximate full adder based on FELIX logicthanks: Preprint Submitted to arXiv

@@ Abstract
Nowadays, much data must be processed and moved between processing and memory units. New technologies and architectures have emerged to improve system performance and overcome the memory bottleneck. The memristor is a technology with both computing and memory capabilities. In-Memory Computing (IMC) can be performed by applying memristors to stateful design methods. The Fast and Energy-Efficient Logic in Memory (FELIX) logic is a stateful implementation logic. The way computations are performed can be changed to improve performance. Approximate design methods can be applied in error-resilient applications. In this paper, an approximate full adder circuit with exact $C_{out}$ and approximate $Sum$ outputs has been proposed using the FELIX design method for IMC in two different implementation approaches. The memristor count in the proposed FELIX-based Approximate Full Adder (FAFA) in the two proposed implementation approaches is improved by 14.28% and 28.57%, energy consumption is improved by 73.735% and 81.754%, and the number of computational steps in both approaches is improved by 66.66% compared to the exact FELIX-based full adder. Two different scenarios are considered for evaluating the FAFA. The results of error analysis and evaluations of the FAFA in three different image processing and average pooling applications confirmed that FAFA has high accuracy and acceptable performance.
@@
Memristor, Approximate Computing, Approximate Full Adder, In-Memory Computing, Image Processing.
@@ 1 Introduction
Recently, numerous efforts have been undertaken to address the Von-Neumann bottleneck or memory wall. This bottleneck arises from the significant energy consumption and delays in transferring data between memory and processing units [1]. Most of the computational energy comes from the energy required to move data between these units because the energy to access memory increases dramatically along the memory hierarchy (from cache to off-chip memory) [2, 3]. Pedram et al. [4] have stated that the energy consumption of a 16-bit adder in 45 nm Complementary Metal-Oxide Semiconductor (CMOS) technology is 3556 times less than the energy needed to access Dynamic Random Access Memory (DRAM). Also, the DRAM access time is 25 times higher than the 32-bit adder delay in CMOS technology [2]. This problem significantly reduces system performance in applications where the volume of data is large and calculations are complex (e.g., machine learning, neural networks, video, and image processing). The physical proximity of memory and processing units was proposed to reduce energy consumption and data movement delay [5]. However, this solution is only able to eliminate data movement partially. On the other hand, reducing the size of MOS Field Effect Transistors (MOSFETs) has faced limitations. These limitations include reduced reliability and gate control, increased subthreshold leakage currents, leakage power, and a cost wall [6, 7, 8]. Consequently, alternative technologies are required to overcome the above problems. Research has been carried out in the past decades, which resulted in a shift towards emerging technologies such as Carbon Nanotube FET (CNFET), Quantum dot Cellular Automata (QCA), Ferroelectric FET (FEFET), nanomagnet logic, and memristors [3, 9].
@@
Adders and multipliers are applied in almost every arithmetic calculation in a computer system [13]. Therefore, because the performance of the systems is influenced by the wide applicability of these units in numerous applications and the high volume of data required for processing, an approximate full adder circuit for IMC has been proposed in this paper. The proposed FELIX-based Approximate Full Adder (FAFA) can be used well in error-resilient applications since the erroneous bit is not propagated from the Least Significant Bits (LSBs) structure to the Most Significant Bits (MSBs). The FAFA has been designed using the FELIX logic method, which is compatible with the structure of crossbar arrays. The key contribution of this paper is proposing a high-speed approximate full adder implementation algorithm (FAFA) with low energy consumption applicable in IMC alongside presenting an approximate truth table with exact $C_{out}$. Optimizing the execution algorithm of the $1^{st}$ implementation approach (FAFA1), following the basic principles of the FELIX method, and presenting a new optimized implementation algorithm (FAFA2) to improve the circuit-level evaluation criteria significantly was the main novelty of the authors in this paper. We calculate circuit evaluation metrics improvements of the FAFA1 and FAFA2 compared to the exact FELIX-based full adder and different exact and approximate IMPLY-based full adders. In order to ensure the output’s accuracy, error analysis metrics in different scenarios are assessed. Then, we apply the FAFA circuit in image processing and average pooling applications and evaluate it according to valid measurements to analyze the quality of images.
@@ 2.1 Memristors
In 1971, Leon Chua proposed the memristor as the fourth fundamental component linking flux and charge [23]. It wasn’t until 2008 that Stanley Williams’ team at HP Labs successfully constructed the first memristor [24]. Following this, researchers have investigated and utilized memristors in various fields. A memristor is a switchable resistor with two electrodes and an insulating active layer between them. Figure 1 shows a memristor [25]. The resistance of a memristor is influenced by the previous voltage or current that has been applied to it. The value of this resistor can change between two values, $R_{on}$ (low resistance state, logic ‘1’) and $R_{off}$ (high resistance state, logic ‘0’), based on the voltage or current applied to it. Suppose the current is applied to the Top Electrode (TE) of the memristor (as shown in Figure 1) or the voltage applied to the memristor is greater than the threshold voltage of the memristor. In that case, the memristor’s resistance decreases and reaches $R_{on}$. On the other hand, when the current is applied to the Bottom Electrode (BE) of the memristor or a voltage below the negative threshold is applied, the resistance increases, causing the memristor to switch to its $R_{off}$ state.
@@ 2.2 FELIX Design Method
Among the memristor-based design methods, one of the inputs is rewritten in the IMPLY method. In the MAGIC method, only the NOR function is implemented in the crossbar array, and all other functions are obtained using NOR, which leads to unnecessary delay overhead. The MRL method is unsuitable for IMC; the FELIX design method is applied in this paper. The FELIX design method enables the activation of IMC and many Boolean functions such as OR, NOT, NOR, Minority (MIN), and NAND in only one cycle and the Majority (MAJ), XOR, and AND functions in two cycles without any additional memristor [16]. While implementing OR in MAGIC requires two NOR cycles, and implementing MAJ and AND requires four cycles in this design method [16]. Like the MAGIC design method, the FELIX design method preserves inputs, and thus, the input data set can be reused for multiple operations [30]. In the FELIX design method, to switch the memristor from $R_{off}$ to $R_{on}$, $V_{pn}>|v_{on}|$ must be established, and $V_{np}>v_{off}$ must be established for switching from $R_{on}$ to $R_{off}$. $|v_{on}|$ and $v_{off}$ are the memristor’s threshold voltages, and $V_{pn}$ ($V_{np}$) is the voltage difference between the $p$ ($n$) and $n$ ($p$) terminals [16]. In FELIX, the execution voltage, $V_{0}$, is a key factor. A specific $V_{0}$ voltage for each logic function should be applied to operate. In other words, the applied voltage determines the type of operation. Figure 2 shows the implementation of 3-input NOR, 3-input NAND, and MIN gates [16]. As shown in Figure 2, these three gates have the same implementation, and the applied voltage value, $V_{0}$, determines the type of operation. These operations require only one cycle. For example, implementing a 3-input NOR gate in a memristor-based crossbar array is described. In the first step, the output memristor (labeled out in Figure 2) is initialized to $R_{on}$. In the next step, $V_{0}$ is applied to the TE (terminal $p$) of the inputs, and the TE of the output memristor is connected to the ground.
@@ 2.3 Memristor-Based Adders
Several methods have been proposed in the literature for designing and implementing full adder cells using memristors. These methods include IMPLY, MAGIC, FELIX, reconfigurable NAND/AND, and MRL. The authors in [21, 31, 32, 33] introduced IMPLY-based exact full adders using serial architecture. Considering the initialization of work memristors, the full adder in [21] needs five memristors. It computes the outputs in 22 computational steps, and the full adder proposed in [31] needs five memristors and 23 computational steps to compute $Sum$ and $C_{out}$. The implementation algorithms for both exact full adders proposed in [32, 33] use three work memristors. Due to this feature and the initialization of three work memristors in a single computational step, which differs from the method used in [21, 31], the outputs of the full adder cells are obtained in 18 [32] and 20 [33] computational steps and stored in the input memristors. In [34], an exact Carry Select Adder (CSA) and an exact Conditional Carry Adder (CCA) are introduced using parallel architecture and with the IMPLY method. The proposed 4-bit CSA and CCA require 44 and 49 memristors, respectively, and take 39 and 41 computational steps, respectively [34]. Compared to serial architecture, parallel architecture has the advantage of being faster. Still, it has disadvantages, such as the need for switches, more memristors, high hardware complexity, and a lack of full compatibility with crossbar arrays [1].
In [16], in addition to introducing the FELIX method and implementing basic logic gates, an exact full adder cell for IMC using XOR and MAJ functions is proposed. In [16], two XOR functions are required to generate $Sum$ output, which takes four cycles and requires two additional memristors. The $C_{out}$ output is also generated after two cycles and requires two additional memristors. It should be mentioned that the initialization of memristors also requires two computational cycles, so the number of cycles increases to eight cycles.
Implementing a full adder using the MAGIC method requires twelve NOR cycles and twelve additional memristors compared to six cycles and four additional memristors in FELIX [35]. The operations related to the exact full adder proposed in [16] and its memristive circuit are illustrated in Table 1 and Figure 3, respectively. In Table 1 and Figure 3, $A_{in}$, $B_{in}$, and $C_{in}$ are the inputs of the full adder, and $W_{1}$ to $W_{4}$ are the four additional memristors needed to perform the addition operation.
@@
Figure 3: Circuit-level design of the FELIX-based full adder cell in [16].
@@
The authors in [1, 22, 37, 38] proposed approximate full adders in the serial architecture using the IMPLY method. The design proposed in [1] requires four memristors and seven computational steps to calculate the outputs, while the best approximate full adder proposed in [22] needs four memristors and eight computational steps. Two serial approximate full adders, SAID1 and SAID2, are introduced using the IMPLY method in [38]. Both of these circuits calculate the outputs of the full adder approximately. The computational accuracy of SAID2 is higher than that of the other circuit designed in [38], but the goal of designing SAID1 was to minimize the circuit complexity. The $Sum$ of the SAID1 is obtained in the second computational step without any work memristor, and its $C_{out}$ is also treated as equivalent to the second input of the full adder, without any calculation. The number of computational steps and the number of memristors required for SAID2 are 6 and 5, respectively [38]. Several approximate adders are presented in [37] using two approximate full adder implementation methodologies. The designs based on the serial architecture that best harmonize with a crossbar array structure are called SINC and SINC+ in [37]. The SINC and SINC+ full adders compute the outputs in 3 and 6 steps, respectively, using 3 and 4 memristors. In the four serial full adders, SAID1, SAID2, SINC, and SINC+, the input memristors are overwritten, and if the input data is needed for other computations, a copy operation must be performed before the algorithm is fully executed.
@@
SAID1 [38] 4 $\frac{4}{8}$ $\frac{2}{8}$ 0.5 0.166 3 2 1 ✓ (IMPLY) Approximate
SAID2 [38] 3 $\frac{3}{8}$ $\frac{2}{8}$ 0.375 0.125 5 6 1 ✓ (IMPLY) Approximate
SINC [37] 6 $\frac{4}{8}$ $\frac{4}{8}$ 0.75 0.25 3 3 1 ✓ (IMPLY) Approximate
SINC+ [37] 4 $\frac{4}{8}$ $\frac{2}{8}$ 0.5 0.166 4 6 1 ✓ (IMPLY) Approximate
@@
[32] 0 0 0 0 0 6 18 1 ✓ (IMPLY) Exact
[33] 0 0 0 0 0 6 20 1 ✓ (IMPLY) Exact
[16] 0 0 0 0 0 7 8 - ✓ (FELIX) Exact
[35] 0 0 0 0 0 15 13 - ✓ (MAGIC) Exact
@@ 3 Proposed Memristive Approximate Full Adder Based On FELIX
In the second approach, the proposed circuit involves five memristors (three input memristors, $A_{in}$, $B_{in}$, and $C_{in}$, and two additional memristors, $W_{1}$ and $W_{2}$). The whole operation is performed in two cycles: one to calculate the MIN function and one to perform the NOT operation. In both implementation approaches, outputs are calculated in three cycles if the initialization of memristors is considered. The implementation algorithms of the FAFA1 and FAFA2 are written in Tables 4 and 5, respectively, and their circuit-level designs are shown in Figures 4(a) and 4(b).
@@ 4.1 Circuit-Level Simulation And Analysis
The LTSPICE simulator verified the correctness of the outputs of all eight possible input states of the FAFA1 and FAFA2. Circuit-level simulation results showed that the proposed algorithm produces outputs corresponding to the proposed truth table in all states. The outputs of the proposed circuit in two input states, $A_{in}B_{in}C_{in}$=“111” and $A_{in}B_{in}C_{in}$=“010”, in the first and second approaches are shown in subfigures 5(a)-5(d).
@@
Table 7 lists the circuit characteristics of the stateful memristive exact and approximate full adders proposed in [1, 21, 22, 32, 33, 37, 38], which are implemented using the IMPLY design method in serial architecture. This table compares only the number of required memristors and the computational steps of the proposed approximate full adders (FAFA1 and FAFA2) with the exact and approximate IMPLY-based full adders. It should be noted that the initialization steps are not included in the computational cycles of the exact and approximate full adders reported in Table 7. Due to the difference in the memristor model used in the circuit level simulation of this paper with the proposed circuits in [1, 21, 22, 32, 33, 37, 38], it is not possible to directly compare the energy consumption of these circuits with each other. The number of memristors applied in the IMPLY-based exact and approximate full adders proposed in [1, 21, 22, 32, 33, 37, 38] is at most 50% less than the number of memristors applied in the proposed circuits here. However, it should be noted that the IMPLY method is destructive, and the input values stored in the input memristors are erased after the implementation algorithm is executed. Therefore, if input data needs to be stored, a copy operation (two NOT operations for each input bit) must be performed using work memristors, significantly increasing the number of computational steps and the number of memristors. The FELIX method is non-destructive, and there is no need to perform a copy operation to preserve the data after executing the proposed circuit implementation algorithm. Proposed approximate full adders have improved the number of computational steps by up to 90% compared to the proposed circuits in [1, 21, 22, 32, 33, 37, 38]. The mentioned percentages are reported without considering the computational steps needed to store the input data in the IMPLY-based exact and approximate full adders [1, 21, 22, 32, 33, 37, 38].
Figure 5: The output waveforms of the proposed full adders: (a) FAFA1: $A_{in}B_{in}C_{in}=``010$", (b) FAFA1: $A_{in}B_{in}C_{in}=``111$", (c) FAFA2: $A_{in}B_{in}C_{in}=``010$", and (d) FAFA2: $A_{in}B_{in}C_{in}=``111$".
Table 7: Exact and approximate full adders’ circuit-level criteria comparison.
@@
Exact IMPLY [21] 5 28.57% 20 -233.33% NC* NC*
Exact IMPLY [32] 6 14.28% 17 -183.33% NC* NC*
Exact IMPLY [33] 6 14.28% 19 -216.66% NC* NC*
SIAFA1 [22] 4 42.85% 7 -16.66% NC* NC*
SIAFA2 [22] 5 28.57% 8 -33.33% NC* NC*
SIAFA3 [22] 4 42.85% 7 -16.66% NC* NC*
SIAFA4 [22] 4 42.85% 7 -16.66% NC* NC*
SAFAN [1] 4 42.85% 6 0% NC* NC*
SAID1 [38] 3 57.14% 2+ 66.66% NC* NC*
SAID2 [38] 5 28.57% 4 33.33% NC* NC*
SINC [37] 3 57.14% 2 66.66% NC* NC*
SINC+ [37] 4 42.85% 5 16.66% NC* NC*
@@
* Not Calculated.
+ The initialization step must be considered because one of the input memristors is directly initialized.
For a comprehensive analysis of the FAFA, the 8-bit approximate Ripple Carry Adder (RCA) has been evaluated in two different scenarios. This approximate adder was divided into two exact and approximate parts in both scenarios. In the exact part of the $1^{st}$ scenario’s structure, the exact full adder was applied to the four MSBs. The FAFA was applied in the LSBs in this scenario. In the $2^{nd}$ scenario, the exact section contains three exact full adders in the MSBs, and the other five are approximate ones (LSBs). In both scenarios, the exact full adder proposed in [16] is used for the MSBs. The $1^{st}$ scenario’s structure is drawn in Figure 6. The memristor count, cycles, and energy consumption (excluding the energy required to initialize the memristors) of the approximate 8-bit RCAs (consisting of FAFA1 and FAFA2) in both scenarios are compared with the exact 8-bit RCA built using the exact full adder proposed in [16] in Table 8.
@@
Table 8: Comparison of the number of memristors, the number of cycles, and the energy consumption of the 8-bit approximate and exact RCAs with the 8-bit exact RCA proposed in [16].
@@
SIAFA1 [22] 19 32.14% 120 -87.5% NC* NC*
SIAFA2 [22] 19 32.14% 128 -100% NC* NC*
SIAFA3 [22] 19 32.14% 120 -87.5% NC* NC*
SIAFA4 [22] 19 32.14% 120 -87.5% NC* NC*
SAFAN [1] 19 32.14% 116 -81.25% NC* NC*
SAID1 [38] 19 32.14% 96 -50% NC* NC*
SAID2 [38] 23 17.86% 112 -75% NC* NC*
SINC [37] 19 32.14% 100 -56.25% NC* NC*
SINC+ [37] 19 32.14% 103 -60.94% NC* NC*
@@
SIAFA1 [22] 19 32.14% 106 -65.62% NC* NC*
SIAFA2 [22] 19 32.14% 116 -81.25% NC* NC*
SIAFA3 [22] 19 32.14% 106 -65.62% NC* NC*
SIAFA4 [22] 19 32.14% 106 -65.62% NC* NC*
SAFAN [1] 19 32.14% 101 -57.81% NC* NC*
SAID1 [38] 19 32.14% 76 -18.75% NC* NC*
SAID2 [38] 24 14.29% 96 -50% NC* NC*
SINC [37] 19 32.14% 81 -26.56% NC* NC*
SINC+ [37] 19 32.14% 84 -31.25% NC* NC*